# Empirical and Sustainability Aspects of Software Engineering Research in the Era of Large Language Models: A Reflection

### v2

@@ Abstract.
conference: ICSE-NIER; 2026; Rio de Janeiro, Brazilccs: Software and its engineering Empirical software validationccs: General and reference Empirical studiesccs: Computing methodologies Artificial intelligenceccs: Social and professional topics Sustainability
@@ 1. Introduction
To understand how our community is tackling the aforementioned challenges, we analyse the empirical research literature published in the research track of the top SE conference, the International Conference on Software Engineering (ICSE), between 2023 and 2025. We identified 177 relevant papers and manually extracted several key pieces of information, including the SE tasks addressed, the types of models used, the experimental protocols, the benchmarks, and other details relevant to empirical studies. We manually analysed this information to establish:
@@
RQ4. Sustainability. What are the costs of LLM-based SE research? Training and deploying LLMs is resource-intensive, with high financial and environmental costs. By examining how sustainability is addressed, we aim to raise awareness of these hidden costs and encourage the community to adopt more transparent, responsible practices that balance innovation, accessibility, and environmental responsibility. We check (i) which experimental costs are reported, (ii) the costs of LLM-based SE research, and (iii) how they are sustained.
@@ 2. Methodology
First, we filtered the initial 692 papers based on the following list of keywords: “LLM”, “language model”, “conversational AI”, “chatbot”, “genAI”, “generative”, “AI-assisted”, “foundation model”, “BERT”, “encoder”, “decoder”, “autoencoder”, “transformer”, and “agent”, retrieving 304 candidates. Among these, we manually selected relevant papers matching the following inclusion criteria: articles that present an empirical experiment involving at least one LLM, either as the main approach or as a baseline. For instance, we exclude papers that focus solely on qualitative insights from LLM user studies. When referring to LLMs, we use the definitions proposed by Fan et al. (Fan et al., 2023), namely encoder-only (e.g., BERT, DeBERTa, RoBERTa), encoder-decoder (e.g., T5, BART), and decoder-only (e.g., GPT, LLaMA, Claude). Our selection followed a three-step process widely adopted in previous surveys (e.g.  (Martin et al., 2016; Hort et al., 2021; Moussa and Sarro, 2022)): (i) We excluded all papers whose titles clearly did not meet our inclusion criteria; (ii) next, we checked abstracts; (iii) finally, we reviewed the full content of articles remaining from the previous two steps and excluded those that did not meet our inclusion criteria. This process resulted in 174 papers selected for data extraction. To mitigate the risks of missing relevant papers, we (i) checked whether we missed any keywords by computing word clouds from the abstracts of the 174 papers (see  (Williams et al., 2025) for full results), finding no additional relevant keywords; (ii) randomly sampled 20% of the 388 papers not selected by our keyword search for each year. This comprised 80 papers (32 for 2023, 28 for 2024, and 20 for 2025), which we manually inspected using the process above. We found only three false negatives and included them in the survey. Thus, we analyse 177 papers in total, with the full list available in our online repository (Williams et al., 2025).
@@ 3. Results & Reflections
We observe a shift in the technologies adopted, from BERT-based models in 2023 to GPT-based models in 2025. While the adoption of more complex and costly technologies is increasing, there does not seem to be a corresponding increase in benchmarking them against non-LLM counterparts. This should be prioritised to ensure the adoption of these technologies remains cost-effective.
Targeted Programming Languages. We examined the prevalence of programming languages (PLs) used in LLM-based empirical studies and found that Java was the most popular language overall, followed by Python, C, C++, and JavaScript. Interestingly, Table 2 shows that Java is becoming less common in studies over the years (from 43.8% of studies in 2023 to 37.9% in 2025), while Python has grown more popular (28.1% to 31.6%). In terms of evaluating across multiple PLs, we see more concerning trends. Overall, only 27.1% of papers evaluate with more than one PL. Notably, only 22.2% of 2025 papers evaluate with more than one PL, compared to the more positive results of 2023 (31.2%) and 2024 (32.7%). This is problematic for replicability, as LLMs have been shown to perform better on some PLs than others (Twist et al., 2025), implying that reported performance on a single PL does not necessarily generalise to other PLs or contexts.
@@
Some papers acknowledged the issue of data contamination without providing strategies to minimise the risk to the validity of the results (Ahmed et al., 2024). In other papers, a common approach to addressing contamination is temporal filtering, which involves collecting data created after the training cutoff date of the evaluated models (Jiang et al., 2025; Wei et al., 2024; Zhang et al., 2025b; Wu et al., 2025). For example, papers use datasets like GitBug-Java (Bouzenia et al., 2025), JLeaks (Wang et al., 2025), DebugBench (Xu et al., 2025), and ClassEval (Zhang et al., 2025b) that were constructed specifically to minimise overlap. Another strategy is benchmark design through obfuscation or perturbation, such as OBFUSEVAL (Zhang et al., 2025a) or DS-1000 (Ouyang et al., 2025), which alters code or phrasing while preserving functionality so that models cannot rely on memorised examples. Data cleaning (Nie et al., 2025), data splitting (Ding et al., 2024), and deduplication are widely applied, with clone detection (Yang et al., 2024) used to eliminate duplicates across training, validation, and test sets. In many cases, experiments are replicated across multiple datasets or with synthetic benchmarks specifically designed to minimise leakage risk. Data contamination is further mitigated by cross-dataset evaluation (Batole et al., 2025) and ablation studies (Ke, 2025; Wang et al., 2025; Guo et al., 2025). By evaluating on multiple benchmarks, including both classic and recently introduced datasets, researchers can assess whether improvements hold beyond data that may be contaminated. Ablation studies, meanwhile, demonstrate that performance gains stem from specific methods or frameworks rather than from data memorisation, since results drop when key components are removed.
@@
While no strategy can fully eliminate the risk of contamination when the training dataset is unavailable, we recommend a combination of temporal filtering, benchmark design consideration, deduplication, cross-dataset validation, and ablation studies to greatly reduce its impact. These measures ensure results reflect model capabilities rather than memorisation and provide the community with practical standards for trustworthy LLM evaluation.
@@
Artefact Availability. In our analysis, we find that 33 out of 177 (18.6%) of empirical studies involving LLMs received an “Artifact Available” badge. This is in stark contrast to the remaining 515 articles published at ICSE, for which 213 badges were awarded (41.4%), thus suggesting that authors conducting research involving LLMs have been less proactive in submitting their contributions to the artefact evaluation tracks. We observe similar trends for “Artifact Reusable” badges (11.9% LLM-based vs. 29.1% other) and “Artifact Functional” badges (7.3% LLM-based vs. 13.6% other). Over the three years examined, we observe a minor increase in the number of studies awarded availability badges. Per year, 6 of the 32 papers (18.75%) from 2023 have these badges, 9 out of 55 papers (16.36%) in 2024, and 18 out of 90 papers (20.00%) in 2025.
@@
Time is the next most frequently discussed cost, being mentioned 68 times. This includes any notion of time reported (e.g., total execution time or time required for a single LLM invocation). Thirty-six publications (20%) reported both the computational environment and time costs, which are necessary to estimate the energy requirements for conducting experiments. This is an improvement over the 10.2% (30 out of 293 publications) reported by Hort et al. (Hort et al., 2023), which considered LLM studies on code-related tasks published up to 2022. However, we found no study that directly reported the energy consumed or $CO_{2}$ emissions.

### v3

@@ Abstract.
conference: 2026 IEEE/ACM 48th International Conference on Software Engineering; April 12–18, 2026; Rio de Janeiro, Brazilccs: Software and its engineering Empirical software validationccs: General and reference Empirical studiesccs: Computing methodologies Artificial intelligenceccs: Social and professional topics Sustainability
@@ 1. Introduction
To understand how our community is tackling the aforementioned challenges, we analyse the empirical research literature published in the research track of the top SE conference, the International Conference on Software Engineering (ICSE), between 2023 and 2025. We identified 177 relevant papers and manually extracted several key pieces of information, including types of models used, experimental protocols, benchmarks, reported costs, and other details relevant to empirical studies. We manually analysed this information to establish:
@@
RQ4. Sustainability. What are the costs of LLM-based SE research? Training and deploying LLMs is resource-intensive, with high financial and environmental costs. By examining how sustainability is addressed, we aim to raise awareness of these hidden costs and encourage the community to adopt more transparent, responsible practices that balance innovation, accessibility, and environmental responsibility. We check (i) which experimental costs are reported, (ii) the costs of LLM-based SE research, and (iii) how they are sustained. We support these findings with a discussion of 57 responses to a questionnaire about experiment costs, which we distributed to authors of the 177 relevant papers.
@@ 2. Methodology
First, we filtered the initial 692 papers based on the following keywords: “LLM”, “language model”, “conversational AI”, “chatbot”, “genAI”, “generative”, “AI-assisted”, “foundation model”, “BERT”, “encoder”, “decoder”, “autoencoder”, “transformer”, and “agent”, retrieving 304 candidates. Among these, we manually selected relevant papers matching the following inclusion criteria: articles that present an empirical experiment involving at least one LLM, either as the main approach or as a baseline. For instance, we exclude papers that focus solely on qualitative insights from LLM user studies. When referring to LLMs, we use the definitions proposed by Fan et al. (Fan et al., 2023), namely encoder-only (e.g., BERT, DeBERTa, RoBERTa), encoder-decoder (e.g., T5, BART), and decoder-only (e.g., GPT, LLaMA, Claude). Our selection followed a three-step process widely adopted in previous surveys (e.g.  (Martin et al., 2016; Hort et al., 2021; Moussa and Sarro, 2022)): (i) We excluded all papers whose titles clearly did not meet our inclusion criteria; (ii) next, we checked abstracts; (iii) finally, we reviewed the full content of articles remaining from the previous two steps and excluded those that did not meet our inclusion criteria. This process resulted in 174 papers selected for data extraction. To mitigate the risks of missing relevant papers, we (i) checked whether we missed any keywords by computing word clouds from the abstracts of the 174 papers (see  (Williams et al., 2025) for full results), finding no additional relevant keywords; (ii) randomly sampled 20% of the 388 papers not selected by our keyword search for each year. This sample comprised 80 papers (32 for 2023, 28 for 2024, and 20 for 2025), which we manually inspected using the process above. We found only three false negatives and included them in the survey. Thus, we analyse 177 papers in total, with the complete list available in our online repository (Williams et al., 2025).
@@ 3. Results & Reflections
We observe a shift in the technologies adopted, from BERT-based models in 2023 to GPT-based models in 2025. While the adoption of more complex and costly technologies increases, there does not seem to be a corresponding increase in benchmarking them against non-LLM counterparts. This should be prioritised to ensure the adoption of these technologies remains cost-effective.
Targeted Programming Languages. We examined the prevalence of programming languages (PLs) used in LLM-based empirical studies and found that Java was the most popular language overall, followed by Python, C, C++, and JavaScript. Interestingly, Table 2 shows that Java is becoming less common in studies over the years (from 43.8% of studies in 2023 to 37.9% in 2025), while Python has grown more popular (28.1% to 31.6%). In terms of evaluating across multiple PLs, we see more concerning trends. Overall, only 27.1% of papers evaluate with more than one PL. Notably, only 22.2% of 2025 papers evaluate with more than one PL, compared to the more positive results of 2023 (31.2%) and 2024 (32.7%). This trend is problematic for replicability, as LLMs have been shown to perform better on some PLs than others (Twist et al., 2025), implying that results reported on one PL will not necessarily generalise to other PLs or contexts.
@@
Some papers acknowledged the issue of data contamination without providing strategies to minimise the risk to the validity of the results (Ahmed et al., 2024). In other papers, a common approach to addressing contamination is temporal filtering, which involves collecting data created after the training cutoff date of the evaluated models (Jiang et al., 2025; Wei et al., 2024; Zhang et al., 2025a; Wu et al., 2025). For example, papers use datasets like GitBug-Java (Bouzenia et al., 2025), JLeaks (Wang et al., 2025), DebugBench (Xu et al., 2025), and ClassEval (Zhang et al., 2025a) that were constructed specifically to minimise overlap. Another strategy is benchmark design through obfuscation or perturbation, such as OBFUSEVAL (Zhang et al., 2025b) or DS-1000 (Ouyang et al., 2025), which alters code or phrasing while preserving functionality so that models cannot rely on memorised examples. Data cleaning (Nie et al., 2025), data splitting (Ding et al., 2024), and deduplication are widely used, with clone detection (Yang et al., 2024) employed to eliminate duplicates across the training, validation, and test sets. In many cases, experiments are replicated across multiple datasets or with synthetic benchmarks specifically designed to minimise leakage risk. Data contamination is further mitigated by cross-dataset evaluation (Batole et al., 2025) and ablation studies (Ke, 2025; Wang et al., 2025; Guo et al., 2025). By evaluating on multiple benchmarks, including both classic and recently introduced datasets, researchers can assess whether improvements hold beyond data that may be contaminated. Ablation studies, meanwhile, demonstrate that performance gains stem from specific methods or frameworks rather than from data memorisation, since results drop when key components are removed.
@@
While no strategy can entirely eliminate the risk of contamination when the training dataset is unavailable, we recommend a combination of temporal filtering, benchmark design consideration, deduplication, cross-dataset validation, and ablation studies to greatly reduce its impact. These measures ensure results reflect model capabilities rather than memorisation and provide the community with practical standards for trustworthy LLM evaluation.
@@
Artefact Availability. In our analysis, we find that 33 out of 177 (18.6%) of empirical studies involving LLMs received an “Artifact Available” badge. This finding is in stark contrast to the remaining 515 articles published at ICSE, for which 213 badges were awarded (41.4%), thus suggesting that authors conducting research involving LLMs have been less proactive in submitting their contributions to the artefact evaluation tracks. We observe similar trends for “Artifact Reusable” badges (11.9% LLM-based vs. 29.1% other) and “Artifact Functional” badges (7.3% LLM-based vs. 13.6% other). Over the three years examined, we observe a minor increase in the number of studies awarded availability badges. Per year, 6 of the 32 papers (18.75%) from 2023 have these badges, 9 out of 55 papers (16.36%) in 2024, and 18 out of 90 papers (20.00%) in 2025.
@@
Time is the next most frequently discussed cost, being mentioned 68 times. This includes any reported notion of time (e.g., total execution time or the time required for a single LLM invocation). Thirty-six publications (20%) reported both the computational environment and time costs, which are necessary to estimate the energy requirements for conducting experiments. This is an improvement over the 10.2% (30 out of 293 publications) reported by Hort et al. (Hort et al., 2023), which considered LLM studies on code-related tasks published up to 2022. However, we found no study that directly reported the energy consumed or $CO_{2}$ emissions.